Determi ne the sample size needed if we wish to construct a confidence interval on the mean of a population when the maximum error allowed is 1/4 standard deviatian , and α 0.01
Answer #1

here from given information margin of error /std deviation =1/4

or σ/E =4

for 99 % CI value of z= 2.576
required sample size n=(zσ/E)2                                         = 107.0